# A Framework for Safe Probabilistic Invariance Verification of Stochastic Dynamical Systems

### v1

@@ 7.1 Discrete-time Systems
In this example we consider the one-dimensional discrete-time system:
@@
The set $\widehat{\mathcal{X}}=\{\,x\mid\widehat{h}(x)\leq 0\,\}$ with $\widehat{h}(x)=x^{2}-2$ is used in solving SDP1-SDP4. The computed lower and upper bounds are summarized in Table 1 and Fig. 1. It is concluded that tighter lower and upper bounds of the liveness probability can be obtained when polynomials of higher degree are used for performing computations. Given that the liveness probability is strictly greater than zero, as stated in Remark 4.6, the use of SDP2 for computing upper bounds on the liveness probability is precluded. This observation aligns with our experimental findings. Instead, upper bounds for the liveness probability can be obtained by solving SDP4, which closely approximate the results obtained from Monte-Carlo methods as the degree of polynomials increases. Concurrently, we note a noteworthy consistency in the nearly identical lower bounds computed from SDP1 and SDP3, with those derived from SDP0 being the most conservative.
Figure 1: Illustration of computed lower and upper bounds (shown in Table 1) of the liveness probability via solving SDP3 and SDP4
@@
In this example we consider the discrete-time Lotka-Volterra model:
$\begin{cases}x(l+1)=rx(l)-ay(l)x(l),\\ y(l+1)=sy(l)+acy(l)x(l),\end{cases}$ (30)
where $r=0.5$, $a=1$, $s=-0.5+d(l)$ with $d(\cdot)\colon\mathbb{N}\rightarrow\mathcal{D}=[-1,1]$ and $c=1$, $\mathcal{X}=\{\,(x,y)^{\top}\mid h(\bm{x})\leq 0\,\}$ $h(\bm{x})=x^{2}+y^{2}-1$, and $\mathcal{X}_{0}=\{\,(x,y)^{\top}\mid g(\bm{x})\leq 0\,\}$ with $g(\bm{x})=(x+0.6)^{2}+(y+0.6)^{2}-0.01$. Besides, we assume that the probability distribution imposed on $\mathcal{D}$ is the uniform distribution. The lower and upper bounds of the liveness probability obtained by Monte Carlo methods are $\epsilon_{1}=0.3651$ and $\epsilon_{2}=0.7974$, respectively. Five trajectories starting from $(-0.6,-0.6)^{\top}$ are visualized in Fig. 3.
The set $\widehat{\mathcal{X}}=\{\,(x,y)^{\top}\mid\widehat{h}(\bm{x})\leq 0\,\}$ with $\widehat{h}(\bm{x})=x^{2}+y^{2}-4$ is used in solving SDP1-SDP4. The computed lower and upper bounds are presented in Table 2. Analogous to Example 7.34, SDP2 cannot be used to compute upper bounds of the liveness probability. Instead, one can obtain upper bounds by solving SDP4, which closely approximates the results obtained from Monte Carlo methods as the degree of polynomials increases. In addition, the lower bounds obtained by solving SDP3 and SDP1 are almost equal and more tighter than those obtained by solving SDP0. Fig. 2 depicts a comparison of the computed $v(\bm{x})$ when employing polynomials of degree 18. The $v(\bm{x})$ obtained from SDP3 is greater than or equal to zero for every $\bm{x}\in\mathcal{X}$ and closely resembles the $v(\bm{x})$ obtained from SDP1. Unlike Op1 and Op3, which impose a non-negative requirement on $v(\bm{x})$ for each $\bm{x}\in\widehat{\mathcal{X}}$, Op4 does not have this requirement and allows $v(\bm{x})$ to be negative for some $\bm{x}\in\widehat{\mathcal{X}}$. This is illustrated in Fig. 2.
@@
$\epsilon_{1}$ 0.1487 0.1865 0.2039 0.2147 0.2203 0.2261
$\epsilon_{2}$ 0.9064 0.8670 0.8501 0.8320 0.8272 0.8171
Op1
$\epsilon_{1}$ 0.1487 0.1865 0.2039 0.2147 0.2203 0.2260
Op0
$\epsilon_{1}$ 0.0000 0.0000 0.0000 0.0000 0.1756 0.1976
Table 2: Computed lower and upper bounds of the liveness probability in Example 7.35 ($d$ denote the degree of unknown polynomials involved in the resulting SDPs)
Figure 3: An illustration of four trajectories for system (30) starting from $(-0.6,-0.6)^{\top}$ (red curve - $\partial\mathcal{X}$; green curve - $\partial\mathcal{X}_{0}$).
In accordance with Lemma 4.15, SDP1 and SDP3 are deemed equivalent, and thus, should yield the same lower bounds. However, a marginal disparity in the computed lower bounds is evident in Table 1 and 2. To the best of our knowledge, the origins of these differences can be attributed to two factors. Firstly, floating-point errors play a role, and secondly, the equivalence of the resulting semi-definite programs SDP3 and SDP1 cannot be guaranteed.
@@ 7.2 Continuous-time Systems
Consider the stochastic differential equation adapted from [16],
$\begin{cases}dX_{1}(t,w)=X_{2}(t,w)dt,\\ dX_{2}(t,w)=-(X_{1}(t,w)+X_{2}(t,w)+0.5X_{1}^{3}(t,w))dt\\ \qquad\qquad\quad+(X_{1}(t,w)+X_{2}(t,w))dW(t,w)\end{cases},$ (31)
@@
In this example, we consider two different initial sets: $\mathcal{X}_{0}=\{\,(x_{1},x_{2})^{\top}\mid g_{1}(\bm{x})<0\,\}$ with $g_{1}(\bm{x})=100(x_{1}+0.4)^{2}+100(x_{2}+0.5)^{2}-1$ and $\mathcal{X}_{0}^{\prime}=\{\,(x_{1},x_{2})^{\top}\mid g_{2}(\bm{x})<0,\}$ with $g_{2}(\bm{x})=10(x_{1}+0.4)^{2}+10(x_{2}+0.5)^{2}-1$. The lower and upper bounds of the liveness probability obtained through Monte Carlo methods are $\epsilon_{1}=0.5338$ and $\epsilon_{2}=0.7101$, respectively, when adopting $\mathcal{X}_{0}$. Conversely, when adopting $\mathcal{X}_{0}^{\prime}$, the lower and upper bounds are $\epsilon_{1}=0.2114$ and $\epsilon_{2}=0.8976$, respectively. Three trajectories starting from $(-0.4,0.5)^{\top}$ are visualized in Fig. 4.
The lower and upper bounds computed by solving SDP5 through SDP7 are summarized in Table 3 and Table 4. The results demonstrate that the lower bounds of the liveness probability obtained from SDP5 and SDP6 are equal, regardless of the initial sets. This aligns with the conclusion stated in Lemma 6.30. Meanwhile, the upper bounds can be computed by solving SDP7, which are close to the ones from Monte Carlo methods. Furthermore, due to the inclusion relationship of $\mathcal{X}_{0}\subset\mathcal{X}_{0}^{\prime}$, the lower bounds using $\mathcal{X}_{0}^{\prime}$ are less than or equal to the lower bounds obtained with $\mathcal{X}_{0}$, while the upper bounds using $\mathcal{X}_{0}^{\prime}$ are greater than or equal to the upper bounds computed using $\mathcal{X}_{0}$. These findings are consistent with our experimental results, as presented in Table 3 and Table 4.
Figure 4: An illustration of four trajectories for system (31) starting from $(-0.4,0.5)^{\top}$ (red curve - $\partial\mathcal{X}$; green curve - $\partial\mathcal{X}_{0}$).
Op6 and Op7
@@
$\epsilon_{1}$ 0.3957 0.4217 0.4590 0.4660 0.4675 0.4682 0.4686
$\epsilon_{2}$ 0.7313 0.7279 0.7233 0.7224 0.7216 0.7213 0.7208
Op5
$\epsilon_{1}$ 0.3957 0.4217 0.4590 0.4660 0.4675 0.4682 0.4686
Table 3: Computed lower and upper bounds of the liveness probability in Example 7.36 using $\mathcal{X}_{0}$
Op6 and Op7
@@
$\epsilon_{1}$ 0.0596 0.0721 0.0873 0.0913 0.0919 0.0923 0.0924
$\epsilon_{2}$ 0.9046 0.9032 0.9011 0.9007 0.9004 0.9002 0.9000
Op5
$\epsilon_{1}$ 0.0596 0.0721 0.0873 0.0913 0.0919 0.0923 0.0924
Table 4: Computed lower and upper bounds of the liveness probability in Example 7.36 using $\mathcal{X}_{0}^{\prime}$
@@
$\begin{cases}dX_{1}(t,w)=&(-0.42X_{1}(t,w)-1.05X_{2}(t,w)\\ &-2.3X_{1}^{2}(t,w)-0.56X_{1}(t,w)X_{2}(t,w)\\ &-X_{1}^{3}(t,w))dt+\sigma X_{1}(t,w)dW(t,w)\\ dX_{2}(t,w)=&(1.98X_{1}(t,w)+X_{1}(t,w)X_{2}(t,w))dt\end{cases},$ (32)
@@
In this example, we consider different values of $\sigma$: $\sigma=0.5$ and $\sigma=2$. When $\sigma=0.5$, the lower bounds and upper bounds of the liveness probability obtained through Monte Carlo methods are $\epsilon_{1}=0.0000$ and $\epsilon_{2}=0.4398$, respectively; when $\sigma=2$, $\epsilon_{1}=0.0000$ and $\epsilon_{2}=0.3580$. For these two scenarios, the lower and upper bounds of the liveness probabilities computed by solving SDP5 to SDP7 are summarized in Table 5 and Table 6, respectively. Additionally, Fig. 5 visualizes three trajectories starting from $(0,0.9)^{\top}$.
Regardless of the values of $\sigma$, the lower bounds obtained from SDP5 and SDP6 are zero. Meanwhile, the upper bounds of the liveness probabilities, computed by solving SDP7, vary for these two values of $\sigma$. It is observed that as $\sigma$ increases, the upper bounds decrease and approach the one obtained from Monte Carlo methods.
Figure 5: An illustration of four trajectories for system (32) starting from $(0,0.9)^{\top}$ with $\sigma=0.5$ in the left and $\sigma=2$ in the right (red curve - $\partial\mathcal{X}$; green curve - $\partial\mathcal{X}_{0}$).
Op6 and Op7
@@
$\epsilon_{1}$ 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
$\epsilon_{2}$ 0.5829 0.5768 0.5709 0.5654 0.5604 0.5574 0.5560
Op5
$\epsilon_{1}$ 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Table 5: Computed lower and upper bounds of the liveness probability in Example 7.37 with $\sigma=0.5$
Op6 and Op7
@@
$\epsilon_{1}$ 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
$\epsilon_{2}$ 0.3731 0.3723 0.3718 0.3715 0.3711 0.3709 0.3707
Op5
$\epsilon_{1}$ 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Table 6: Computed lower and upper bounds of the liveness probability in Example 7.37 with $\sigma=2$
@@ 8 Conclusion
This paper introduced several optimizations aimed at addressing the safe probabilistic invariance verification problem in both stochastic discrete-time and continuous-time systems. Specifically, the goal is to compute lower and upper bounds of the liveness probability concerning a safe set and an initial set of states. These optimizations were constructed by adapting either nonnegative supermartingale barrier certificates or equations that characterize exact reach-avoid probabilities. The effectiveness and comparisons of these optimizations were thoroughly evaluated through both theoretical and numerical analyses.
Appendix
8.1 The proof of Lemma 6.30
Proof 8.38.
@@
$\mathbb{E}[v(\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w}))]\leq v(\bm{x}_{0})<-\delta,\forall t\in\mathbb{R}_{\geq 0}.$ (33)
Also since $v(\bm{x})\geq 1_{\partial\mathcal{X}}(\bm{x})+\widehat{\mathcal{A}}u(\bm{x}),\forall\bm{x}\in\overline{\mathcal{X}}$, we have that
$v(\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w}))\geq\mathcal{A}u(\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w})),\forall t\in\mathbb{R}_{\geq 0}.$ (34)
Thus, we have
@@
According to Theorem 5.20 and (33), we have that
@@
$\frac{\mathbb{E}[u(\widehat{\bm{X}}^{\bm{x}_{0}}(\tau,\bm{w}))]-u(\bm{x}_{0})}{\tau}<-\delta,\forall\tau\in\mathbb{R}_{>0}.$ (35)
We have that
$\lim_{\tau\rightarrow\infty}\frac{\mathbb{E}[u(\widehat{\bm{X}}^{\bm{x}_{0}}(\tau,\bm{w}))]-u(\bm{x}_{0})}{\tau}=0,$
@@
which contradicts (35). Therefore, we can conclude that if $v(\bm{x})$ satisfies (28), $v(\bm{x})\geq 0$ holds for $\bm{x}\in\mathcal{X}$. Thus, constraints (26) and (28) are equivalent.
8.2 Semi-definite Programming Implementation

### v2

@@ 7.1 Discrete-time Systems
Consider the one-dimensional discrete-time system:
@@
The set $\widehat{\mathcal{X}}=\{\,x\mid\widehat{h}(x)\leq 0\,\}$ with $\widehat{h}(x)=x^{2}-2.25$ is used in solving SDP1-SDP4. The computed lower and upper bounds are summarized in Table 1. It is concluded that tighter lower and upper bounds of the liveness probability can be obtained when polynomials of higher degree are used for performing computations. Given that the liveness probability is strictly greater than zero, as stated in Remark 4.8, the use of SDP2 for computing upper bounds on the liveness probability is precluded. This observation aligns with our experimental findings. Instead, upper bounds for the liveness probability can be obtained by solving SDP4, which closely approximate the results obtained from Monte-Carlo methods as the degree of polynomials increases. Concurrently, we note a noteworthy consistency in the identical lower bounds computed from SDP1 and SDP3, with those derived from SDP0 being the most conservative. The identical lower bounds computed from SDP1 and SDP3 also justify the validity of the conclusion in Proposition 4.17.
@@
Consider the discrete-time Lotka-Volterra model:
(30) $\begin{cases}x(l+1)=rx(l)-ay(l)x(l),\\ y(l+1)=sy(l)+acy(l)x(l),\end{cases}$
where $r=0.5$, $a=1$, $s=-0.5+d(l)$ with $d(\cdot)\colon\mathbb{N}\rightarrow\mathcal{D}=[-1,1]$ and $c=1$, $\mathcal{X}=\{\,(x,y)^{\top}\mid h(\bm{x})\leq 0\,\}$ with $h(\bm{x})=x^{2}+y^{2}-1$, and $\mathcal{X}_{0}=\{\,(x,y)^{\top}\mid g(\bm{x})\leq 0\,\}$ with $g(\bm{x})=(x+0.6)^{2}+(y+0.6)^{2}-0.01$. Besides, we assume that the probability distribution imposed on $\mathcal{D}$ is the uniform distribution.
The set $\widehat{\mathcal{X}}=\{\,(x,y)^{\top}\mid\widehat{h}(\bm{x})\leq 0\,\}$ with $\widehat{h}(\bm{x})=x^{2}+y^{2}-4$ is used in solving SDP1-SDP4. The lower and upper bounds of the liveness probabilities obtained by Monte Carlo methods are $\epsilon_{1}=0.3651$ and $\epsilon_{2}=0.7974$, respectively. Five trajectories starting from $(-0.6,-0.6)^{\top}$ are visualized in Fig. 2(a). The computed lower and upper bounds are presented in Table 2. Analogous to Example 7.1, SDP2 cannot be used to compute upper bounds of the liveness probabilities. Instead, upper bounds can be obtained by solving SDP4, which closely approximates the results obtained from Monte Carlo methods as the degree of polynomials increases. In addition, the lower bounds obtained by solving SDP3 and SDP1 are almost equal and tighter than those obtained by solving SDP0. Fig. 1 depicts a comparison of the computed $v(\bm{x})$ when employing polynomials of degree 18. The $v(\bm{x})$ obtained from SDP3 is greater than or equal to zero for every $\bm{x}\in\mathcal{X}$ and closely resembles the $v(\bm{x})$ obtained from SDP1. Unlike Op1 and Op3, which impose a non-negative requirement on $v(\bm{x})$ for each $\bm{x}\in\widehat{\mathcal{X}}$, Op4 does not have this requirement and allows $v(\bm{x})$ to be negative for some $\bm{x}\in\widehat{\mathcal{X}}$. This is illustrated in Fig. 1.
@@
$\epsilon_{1}$ by Op3 0.1487 0.1865 0.2039 0.2147 0.2203 0.2261
$\epsilon_{2}$ by Op4 0.9064 0.8670 0.8501 0.8320 0.8272 0.8171
$\epsilon_{1}$ by Op1 0.1487 0.1865 0.2039 0.2147 0.2203 0.2260
$\epsilon_{1}$ by Op0 0.0000 0.0000 0.0000 0.0000 0.1756 0.1976
Table 2: Computed lower and upper bounds of the liveness probabilities in Example 7.2 ($d$ denote the degree of unknown polynomials involved in the resulting SDPs)
(a) Trajectories for system (30) starting from $(-0.6,-0.6)^{\top}$.
(b) Trajectories for system (31) starting from $(-0.4,0.5)^{\top}$.
(c) Trajectories for system (32) starting from $(0,0.9)^{\top}$ with $\sigma=0.5$ in the left and $\sigma=2$ in the right.
Figure 2: Illustrations of trajectories for systems (red curve - $\partial\mathcal{X}$; green curve - $\partial\mathcal{X}_{0}$).
@@ 7.2 Continuous-time Systems
Consider the stochastic differential equation that is adapted from [17],
(31) $\begin{cases}dX_{1}(t,w)=X_{2}(t,w)dt,\\ dX_{2}(t,w)=-(X_{1}(t,w)+X_{2}(t,w)+0.5X_{1}^{3}(t,w))dt\\ \qquad\qquad\quad+(X_{1}(t,w)+X_{2}(t,w))dW(t,w)\end{cases},$
@@
In this example, we consider two different initial sets: $\mathcal{X}_{0}=\{\,(x_{1},x_{2})^{\top}\mid g_{1}(\bm{x})<0\,\}$ with $g_{1}(\bm{x})=100(x_{1}+0.4)^{2}+100(x_{2}+0.5)^{2}-1$ and $\mathcal{X}_{0}^{\prime}=\{\,(x_{1},x_{2})^{\top}\mid g_{2}(\bm{x})<0\}$ with $g_{2}(\bm{x})=10(x_{1}+0.4)^{2}+10(x_{2}+0.5)^{2}-1$. The lower and upper bounds of the liveness probabilities obtained through Monte Carlo methods are $\epsilon_{1}=0.5338$ and $\epsilon_{2}=0.7101$, respectively, when adopting $\mathcal{X}_{0}$. Conversely, when adopting $\mathcal{X}_{0}^{\prime}$, the lower and upper bounds are $\epsilon_{1}=0.2114$ and $\epsilon_{2}=0.8976$, respectively. Three trajectories starting from $(-0.4,0.5)^{\top}$ are visualized in Fig. 2(b).
The lower and upper bounds computed by solving SDP5 through SDP7 are summarized in Table 3 and Table 4. The results demonstrate that the lower bounds of the liveness probabilities obtained from SDP5 and SDP6 are equal, regardless of the initial sets. This aligns with the conclusion stated in Proposition 6.6. Meanwhile, the upper bounds can be computed by solving SDP7, which are close to the ones from Monte Carlo methods. Furthermore, due to the inclusion relationship of $\mathcal{X}_{0}\subset\mathcal{X}_{0}^{\prime}$, the lower bounds using $\mathcal{X}_{0}^{\prime}$ are less than or equal to the lower bounds obtained with $\mathcal{X}_{0}$, while the upper bounds using $\mathcal{X}_{0}^{\prime}$ are greater than or equal to the upper bounds computed using $\mathcal{X}_{0}$. These findings are consistent with our experimental results, as presented in Table 3 and Table 4.
@@
$\epsilon_{1}$ by Op6 0.3957 0.4217 0.4590 0.4660 0.4675 0.4682 0.4686
$\epsilon_{2}$ by Op7 0.7313 0.7279 0.7233 0.7224 0.7216 0.7213 0.7208
$\epsilon_{1}$ by Op5 0.3957 0.4217 0.4590 0.4660 0.4675 0.4682 0.4686
Table 3: Computed lower and upper bounds of the liveness probabilities in Example 7.3 using $\mathcal{X}_{0}$ ($d$ denotes the degree of unknown polynomials involved in the SDPs)
@@
$\epsilon_{1}$ by Op6 0.0596 0.0721 0.0873 0.0913 0.0919 0.0923 0.0924
$\epsilon_{2}$ by Op7 0.9046 0.9032 0.9011 0.9007 0.9004 0.9002 0.9000
$\epsilon_{1}$ by Op5 0.0596 0.0721 0.0873 0.0913 0.0919 0.0923 0.0924
Table 4: Computed lower and upper bounds of the liveness probabilities in Example 7.3 using $\mathcal{X}_{0}^{\prime}$ ($d$ denotes the degree of unknown polynomials involved in the SDPs)
@@
(32) $\begin{cases}dX_{1}(t,w)=&\big{(}-0.42X_{1}(t,w)-1.05X_{2}(t,w)-2.3X_{1}^{2}(t,w)\\ &-0.56X_{1}(t,w)X_{2}(t,w)-X_{1}^{3}(t,w)\big{)}dt+\sigma X_{1}(t,w)dW(t,w)\\ dX_{2}(t,w)=&(1.98X_{1}(t,w)+X_{1}(t,w)X_{2}(t,w))dt\end{cases},$
@@
In this example, we consider different values of $\sigma$: $\sigma=0.5$ and $\sigma=2$. When $\sigma=0.5$, the lower and upper bounds of the liveness probabilities obtained through Monte Carlo methods are $\epsilon_{1}=0.0000$ and $\epsilon_{2}=0.4398$, respectively; when $\sigma=2$, $\epsilon_{1}=0.0000$ and $\epsilon_{2}=0.3580$. For these two scenarios, the lower and upper bounds of the liveness probabilities computed by solving SDP5 to SDP7 are summarized in Table 5 and Table 6, respectively. Additionally, Fig. 2(c) visualizes three trajectories starting from $(0,0.9)^{\top}$.
Regardless of the values of $\sigma$, the lower bounds obtained from SDP5 and SDP6 are zero. Meanwhile, the upper bounds of the liveness probabilities, computed by solving SDP7, vary for these two values of $\sigma$. It is observed that as the degree $d$ of polynomials used for computations increases, the upper bounds decrease and approach the one obtained from Monte Carlo methods.
@@
$\epsilon_{1}$ by Op6 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
$\epsilon_{2}$ by Op7 0.5829 0.5768 0.5709 0.5654 0.5604 0.5574 0.5560
$\epsilon_{1}$ by Op5 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Table 5: Computed lower and upper bounds of the liveness probabilities in Example 7.4 with $\sigma=0.5$ ($d$ denotes the degree of unknown polynomials involved in the SDPs)
@@
$\epsilon_{1}$ by Op6 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
$\epsilon_{2}$ by Op 7 0.3731 0.3723 0.3718 0.3715 0.3711 0.3709 0.3707
$\epsilon_{1}$ by Op5 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Table 6: Computed lower and upper bounds of the liveness probabilities in Example 7.4 with $\sigma=2$ ($d$ denotes the degree of unknown polynomials involved in the SDPs)
Example 7.5.
Consider a 3D VanderPol oscillator adapted from [12],
$\begin{cases}dX_{1}(t,w)=&-2X_{2}(t,w)dt+X_{1}(t,w)dW(t,w)\\ dX_{2}(t,w)=&\big{(}0.8X_{1}(t,w)-2.1X_{2}(t,w)+X_{3}(t,w)+10X_{1}^{2}(t,w)X_{2}(t,w)\big{)}dt\\ &+X_{2}(t,w)dW(t,w)\\ dX_{3}(t,w)=&\big{(}-X_{3}(t,w)+X_{3}^{3}(t,w)\big{)}dt+X_{3}(t,w)dW(t,w)\\ \end{cases},$
where the safe set is $\mathcal{X}=\{\,(x_{1},x_{2},x_{3})^{\top}\mid h(\bm{x})<0\,\}$ with $h(\bm{x})=x_{1}^{2}+x_{2}^{2}+x_{3}^{3}-1$ and the initial set is $\mathcal{X}_{0}=\{\,(x_{1},x_{2},x_{3})^{\top}\mid g(\bm{x})<0\,\}$ with $g(\bm{x})=(x_{1}-0.3)^{2}+(x_{2}+0.2)^{2}+(x_{3}-0.2)^{2}-0.001$.
In this example, the lower bound and upper bound of the liveness probabilities obtained through Monte Carlo methods are $\epsilon_{1}=0.7363$ and $\epsilon_{2}=0.8150$, respectively. The lower and upper bounds of the liveness probabilities computed by solving SDP5 to SDP7 are summarized in Table 7. It is observed that as the degree $d$ of polynomials used for computations increases, the lower and upper bounds approach the ones obtained from Monte Carlo methods.
d 4 6 8 10 12 14 16
$\epsilon_{1}$ by Op6 0.5810 0.6861 0.7027 0.7112 0.7162 0.7178 0.7188
$\epsilon_{2}$ by Op7 0.9598 0.8496 0.8336 0.8266 0.8238 0.8223 0.8218
$\epsilon_{1}$ by Op5 0.5810 0.6861 0.7027 0.7112 0.7162 0.7178 0.7188
Table 7: Computed lower and upper bounds of the liveness probabilities in Example 7.5 ($d$ denotes the degree of unknown polynomials involved in the SDPs)
@@ 8 Conclusion
This paper introduced several optimizations aimed at addressing the safe probabilistic invariance verification problem in both stochastic discrete-time and continuous-time systems. Specifically, the goal is to compute lower and upper bounds of the liveness probabilities concerning a safe set and an initial set of states. These optimizations were constructed via either using the Doob’s nonnegative supermartingale inequality-based method or relaxing the equations that characterize exact reach-avoid probabilities. The effectiveness and comparisons of these optimizations were thoroughly evaluated through both theoretical and numerical analyses.
Appendix A The proof of Proposition 6.6
Proof A.1.
@@
(33) $\mathbb{E}[v(\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w}))]\leq v(\bm{x}_{0})<-\delta,\forall t\in\mathbb{R}_{\geq 0}.$
Also since $v(\bm{x})\geq 1_{\partial\mathcal{X}}(\bm{x})+\widehat{\mathcal{A}}u(\bm{x}),\forall\bm{x}\in\overline{\mathcal{X}}$, we have that $v(\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w}))\geq\mathcal{A}u(\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w})),\forall t\in\mathbb{R}_{\geq 0}$.
@@
According to Lemma 5.2 and (33), we have that
@@
(34) $\frac{\mathbb{E}[u(\widehat{\bm{X}}^{\bm{x}_{0}}(\tau,\bm{w}))]-u(\bm{x}_{0})}{\tau}<-\delta,\forall\tau\in\mathbb{R}_{>0}.$
@@
We have that $\lim_{\tau\rightarrow\infty}\frac{\mathbb{E}[u(\widehat{\bm{X}}^{\bm{x}_{0}}(\tau,\bm{w}))]-u(\bm{x}_{0})}{\tau}=0$, which contradicts (34). Therefore, we can conclude that if $v(\bm{x})$ satisfies (28), $v(\bm{x})\geq 0$ holds for $\bm{x}\in\mathcal{X}$. Thus, constraints (26) and (28) are equivalent.
Appendix B Optimizations for Computing Lower and Upper Bounds